Calculate [OH -] and pH for each of the following solutions.



(a) 0.0092 M CsOH

[OH-] =  M pH =




(b) 0.0335 g of RbOH in 520.0 mL of solution

[OH -] =  M pH =




(c) 62.6 mL of 0.00552 M Ca(OH)2 diluted to 900 mL

[OH -] =  M pH =




(d) A solution formed by mixing 64.0 mL of 0.000150 M Ca(OH)2 with 89.0 mL of 4.3 x 10-3 M CsOH

[OH -] =  M pH =
Answer #1

Solution: AP) lu a given question, Molarity of 65001 solE 0.0092.74 CSOH is a strong base ie (oh discount iu sol 100%. CS OFrom ci) molarity of Solwicy G. ROOH = no. of moles of Rboh X 1000 Volume of soll (ver) 10003269 xlogo 520 0.00062077 = is stAfter ditution no of to Total volume of solº = 900+62:6 So Now Y lors = 962.6 ml Molarity of cpH) soľ - no. of moles Toral vod) buren First sol, vou of ocool50m (atH) a sol = 64 mL no. of moles of calo H), salu sol = Molarity & vol of sol looo = 0.01000) Molarity of sol - Total no of mores . Total wo (** of 500m = 0·01 Yloud 153 00655 m le → LalOH) is 0H) Both strong You